OpenStudy (anonymous):

subtract plse show me how (7/10x3 - 3/10x - 1/8) - (-1/10x3 + 3/10x - 1/8) =

OpenStudy (anonymous):

the minus sign basically negates all the signs of the stuff in the second set of parens to get \[\frac{7}{10}x^3 - \frac{3}{10}x - \frac{1}{8} + \frac{1}{10}x^3 - \frac{3}{10}x + \frac{1}{8}\] Then you combine like terms to get\[\frac{8}{10}x^3 - \frac{6}{10}x\]which simplifies to\[\frac{4}{5}x^3 - \frac{3}{5}x\]

OpenStudy (anonymous):

thanks for the clarification greatly appreciated